### v1

@@ 6.1 A Universal $\mathcal{O}(1/\sqrt[4]{k})$ Rate for General Projections Onto Convex Sets
Projections Onto Convex Sets (POCS) is a classical method that iteratively projects onto closed convex sets to find a point in their intersection (Gubin et al., 1967; Boyd et al., 2003). Formally,
@@
       $\mathbf{w}_{t}\leftarrow{\Pi}_{\tau(t)}(\mathbf{w}_{t-1})\triangleq\operatorname*{argmin}_{\mathbf{w}\in\mathcal{C}_{\tau(t)}}\left\|{\mathbf{w}-\mathbf{w}_{t-1}}\right\|$
@@
Consider $T$ arbitrary (nonempty) closed convex sets $\mathcal{C}_{1},\dots,\mathcal{C}_{T}$, initial point $\mathbf{w}_{0}\in\mathbb{R}^{d}$, and ordering $\tau$. Define $f_{m}(\mathbf{w})=\frac{1}{2}\left\|{\mathbf{w}-{\Pi}_{m}(\mathbf{w})}\right\|^{2},\forall m\in\left[{T}\right]$. Then,
@@
    The POCS update is equivalent to an SGD step: $\mathbf{w}_{t}={\Pi}_{\tau(t)}(\mathbf{w}_{t-1})=\mathbf{w}_{t-1}-\nabla_{\mathbf{w}}f_{\tau(t)}(\mathbf{w}_{t-1})$.
@@
Theorem 14 (Universal POCS Rate).
Consider the same conditions of Reduction 3 and assume a nonempty set intersection $\mathcal{C}_{\star}=\bigcap_{m=1}^{T}\mathcal{C}_{m}\neq\varnothing$. Then, under a random ordering with or without replacement, the expected “residual” of Scheme 4 after $\forall k\geq 1$ iterations (without replacement: $k\in\left[{T}\right]$) is bounded as,
$\mathbb{E}_{\tau}\Big{[}\frac{1}{2T}{\sum}_{m=1}^{T}\left\|{\mathbf{w}_{k}-{\Pi}_{m}(\mathbf{w}_{k})}\right\|^{2}\Big{]}=\mathbb{E}_{\tau}\Big{[}\frac{1}{2T}{\sum}_{m=1}^{T}\mathrm{dist}^{2}(\mathbf{w}_{k},\mathcal{C}_{m})\Big{]}\leq\frac{7}{\sqrt[4]{k}}\min_{\mathbf{w}\in\mathcal{C}_{\star}}\left\|{\mathbf{w}_{0}-\mathbf{w}}\right\|^{2}\,.$
To the best of our knowledge, the above is the first universal rate in the POCS literature, independent of problem parameters such as regularity or complexity measure, as demonstrated in the next Section 6.2. Universal rates are only achievable when analyzing individual distances, i.e., $f_{m}(w)=\left\|{\mathbf{w}-{\Pi}_{m}(\mathbf{w})}\right\|^{2}=\mathrm{dist}^{2}(\mathbf{w},\mathcal{C}_{m})$, rather than the distance to the intersection, i.e., $\mathrm{dist}^{2}(\mathbf{w},\mathcal{C}_{\star})$. In machine learning, the squared distance from individual sets is linked to important losses like MSE in regression or squared hinge loss in classification (Evron et al., 2022, 2023), naturally leading us to our next continual learning model.
@@ 6.2 A Universal $\mathcal{O}(1/\sqrt[4]{k})$ Rate for Random Orderings in Continual Linear Classification
Regularization methods are commonly used to prevent forgetting in continual learning (see Kirkpatrick et al., 2017). Evron et al. (2023) studied a regularized linear model for continual classification. They considered $T\geq 2$ jointly separable, binary classification tasks, defined by datasets $S_{1},\dots,S_{T}$ consisting of vectors $\mathbf{x}\in\mathbb{R}^{d}$ and their labels $y\in\left\{-1,+1\right\}$. They proved that a weakly-regularized scheme implicitly applies sequential max-margin projections. That is, in the limit as $\lambda\to 0$, the iterates of the two following schemes align in direction, enabling the study of continual classification through the lens of projection algorithms.
@@
   For each task $t=1,\dots,k$:
@@
   For each task $t=1,\dots,k$:
       $\mathbf{w}_{t}\leftarrow{\Pi}_{\tau(t)}(\mathbf{w}_{t-1})\triangleq\operatorname*{argmin}_{\mathbf{w}\in\mathcal{C}_{\tau(t)}}\left\|{\mathbf{w}-\mathbf{w}_{t-1}}\right\|$
       where $\mathcal{C}_{m}\!\triangleq\!\left\{\mathbf{w}\in\mathbb{R}^{d}\mid y\mathbf{w}^{\top}\mathbf{x}\geq 1,~{}\forall(\mathbf{x},y)\!\in\!S_{m}\right\}$
They studied forgetting under several orderings, using an equivalent of our Definition 3:
$F_{\tau}(k)=\frac{1}{k}\sum_{t=1}^{k}\big{(}{\mathcal{L}_{\tau(t)}(\mathbf{w}_{k})-\mathcal{L}_{\tau(t)}(\mathbf{w}_{t})}\big{)}\leq\frac{R^{2}}{2k}\sum_{t=1}^{k}\left\|{\mathbf{w}_{k-1}-{\Pi}_{\tau(t)}(\mathbf{w}_{k-1})}\right\|^{2}\,.$
Our POCS rate (Theorem 14) combined with SGD stability arguments give the following.
Theorem 15.
Under a random ordering, with or without replacement, over $T$ jointly separable tasks, the expected forgetting of the weakly-regularized Scheme 5 (at $\lambda\to 0$) after $k\geq 1$ iterations is bounded as
As shown in Table 3, our rate is universal while the previous one is highly dependent on $\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}$, often thought of as the “complexity” of classification problems. For example, after $k=4T\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}$ iterations, it gives a $e^{-1}$ (normalized) forgetting rate while our rate gives a potentially much smaller $\frac{5}{T^{1/4}\sqrt{\left\|\mathbf{w}_{\star}\right\|R}}$.
@@
Evron et al. (2023) $\exp\Big{(}{-\frac{k}{4T\|{\mathbf{w}_{\star}}\|^{2}R^{2}}}\Big{)}$ $\frac{T^{2}}{\sqrt{k}}\wedge\exp\Big{(}{-\frac{k}{16T^{2}\|{\mathbf{w}_{\star}}\|^{2}R^{2}}}\Big{)}$
@@ 7 Discussion
Our work established a reduction from continual linear regression to the (block) Kaczmarz method and then to “stepwise-optimal” SGD. This enabled the development of analytic tools for last-iterate SGD schemes, leading to significantly improved and even universal rates for random orderings in continual learning and the Kaczmarz method. Our main results are summarized in Tables 1 and 2.
Much of the related work has been covered throughout the paper. A further discussion of related work can be found in Appendix A. Here, we briefly highlight additional aspects of our work.
@@
Many popular continual benchmarks in deep learning implicitly assume a random ordering, such as the permuted MNIST benchmark (Kirkpatrick et al., 2017). In this paper, we showed that in sufficiently long task sequences, random ordering is enough to prevent catastrophic forgetting, and the training loss goes to zero, even in the worst case. In accordance with our results, Lesort et al. (2023) examined a continual learning random benchmark—in which a subset of classes is randomly sampled in each task—and observed that forgetting diminishes as the number of sampled tasks increases, even when training with standard SGD (without any modifications to mitigate forgetting). This suggests that random orderings may contaminate continual learning benchmarks, making it harder to isolate the algorithmic effects being tested. Furthermore, real-world tasks often change gradually, not adhering to random orderings. Such “gradually evolving” datasets might be more challenging and perhaps more relevant as continual benchmarks.
@@
In Section 3.1 we revisited known connections between continual regression and the Kaczmarz method (Evron et al., 2022). We broadened this connection in Section 3.2, bridging the block Kaczmarz method and “stepwise-optimal” SGD, thus applying our novel SGD bounds to the Kaczmarz method. Using Kaczmarz terminology, given a system $\mathbf{A}\mathbf{x}=\mathbf{b}$ consisting of $T$ blocks of an average rank $\bar{r}$ where $\mathbf{A}_{m}\in\mathbb{R}^{{n_{m}}\times d},\mathbf{b}_{m}\in\mathbb{R}^{n_{m}}$, our rates from Section 4 can be summarized as $\mathbb{E}_{\tau}\big{[}\frac{1}{2T}\sum_{m=1}^{T}\big{\|}{\mathbf{A}_{m}\mathbf{x}_{k}-\mathbf{b}_{m}}\big{\|}^{2}\big{]}=\mathcal{O}\big{(}{\min\big{(}{k^{-1/4},\frac{1}{k}\sqrt{d-\bar{r}},\frac{1}{k}\sqrt{T\bar{r}}}\big{)}}\big{)}$ for random orderings with replacement and $\mathcal{O}\big{(}{\min\big{(}{k^{-1/4},\frac{1}{k}\left({d-\bar{r}}\right)}\big{)}}\big{)}$ without replacement. Note that we bounded the loss, rather than the “error” $\left\|{\mathbf{w}_{k}-\mathbf{w}_{\star}}\right\|^{2}$, thus enabling the derivation of rates independent of quantities like the condition number that can make convergence arbitrarily slow.
@@
The seminal work of Strohmer and Vershynin (2009) proposed a Kaczmarz method where rows are sampled with probability proportional to their squared norm. Our approach accommodates non-uniform sampling, including the norm-based one, which leads to a tighter version of Theorem 8, replacing the dependence on the maximum row norm $R$ with the average one. In the block version, both uniform and non-uniform variants exist (Needell and Tropp, 2014; Gower and Richtárik, 2015). As in rank-$1$, our approach should yield better bounds when weighting samples proportional to block norms, again improving the dependence on the maximum norm to the average one.
@@
We narrowed the gap between existing lower and upper worst-case bounds for random orderings in continual linear regression (see Table 1). However, a considerable gap remains between $\Omega(1/k)$ and $\mathcal{O}({1/k^{1/4}})$. Generally, we conjecture that the last-iterate SGD rates can be improved beyond those in Theorem 10, and that Theorem 13 can be extended to the multi-epoch setup. Following our reductions (Section 3), improved rates for “stepwise-optimal” SGD rates would immediately refine the bounds for continual linear regression and classification.
Acknowledgments.
@@ Appendix A Related Work
For the general (non-realizable) smooth stochastic setup, the recent work of Liu and Zhou (2024) was the first (and only, to our knowledge) to provide upper bounds on the convergence rate of the last SGD iterate. While their bounds are applicable in the realizable setting, they require non-constant step sizes to obtain non-trivial convergence, and are therefore not useful for our purposes (see Table 2). Our analysis technique in Section 5.1 borrows from the work of Shamir and Zhang (2013, also mentioned in Table 2) which, in fact, belongs to the comparatively-richer line of work on the non-smooth setting (Shamir and Zhang, 2013; Jain et al., 2019; Zamani and Glineur, 2023; Liu and Zhou, 2024). Notably, SGD in a stochastic non-realizable (either smooth or non-smooth) setup requires uniformly bounded noise assumptions, and generally cannot accommodate a constant step size independent of the optimization horizon.
@@
Our analysis for SGD without-replacement is related to a long line of work primarily focused on the average iterate convergence rates (e.g., Recht and Ré, 2012b; Nagaraj et al., 2019; Safran and Shamir, 2020; Rajput et al., 2020; Mishchenko et al., 2020; Cha et al., 2023; Cai et al., 2023). For the non-strongly convex case, near-optimal bounds (for the average iterate) have been established for the general smooth case (Nagaraj et al., 2019; Mishchenko et al., 2020). In a subsequent work, Cai et al. (2023) refined the dependence on problem parameters for the smooth realizable case (among others). Guarantees for the last iterate have only been established recently by Cai and Diakonikolas (2025). However, their bounds decay with the number of epochs rather than the number of iterations and apply only to non-constant step sizes, making them inapplicable to our setting. Specifically, in a realizable $\beta$-smooth setup, after $J$ without-replacement SGD epochs over a finite sum of size $n$, Mishchenko et al. (2020); Cai et al. (2023) obtained an $O(\beta/J)$ bound for the average iterate with step size $\eta=1/(\beta n)$; and Cai and Diakonikolas (2025) derived a similar bound for the last iterate up to logarithmic factors.
@@
Another line of work related to ours studies algorithmic stability (Bousquet and Elisseeff, 2002; Shalev-Shwartz et al., 2010) of gradient methods, which is the main technique we use in the proof of Theorem 13. Our approach is similar in nature to that of Nagaraj et al. (2019); Sherman et al. (2021); Koren et al. (2022) and primarily builds on Sherman et al. (2021), who were the first to formally introduce the notion of without-replacement stability. For with-replacement SGD, Hardt et al. (2016) discussed its algorithmic stability under smooth loss functions. Later, Lei and Ying (2020), improved this bound in the realizable loss case. The case we consider—i.e., the stability of without-replacement SGD under smooth and realizable loss functions—is not covered in the existing literature.
@@
Our results in Section 4 establish universal bounds for random orderings, both with and without replacement. Both the with- and without-replacement variants converge linearly towards the minimum-norm solution $\mathbf{w}_{\star}$ (Gower and Richtárik, 2015; Han and Xie, 2024), but as we explained in Section 7, the rates can be arbitrarily slow. Recht and Ré (2012a) formulated a noncommutative analog of the arithmetic-geometric mean inequality that, if true, could have shown that without-replacement orderings lead to faster loss convergence than with-replacement orderings in Kaczmarz methods, and consequently in continual linear regression. Years later, Lai and Lim (2020) proved that this inequality does not hold in general (see also De Sa, 2020). Moreover, as in other areas, empirical studies found that row shuffling followed by cyclic orderings performs as well as i.i.d. orderings (Oswald and Zhou, 2015). This naturally connects to interesting observations and open questions regarding various forms of shuffled SGD (Bottou, 2009; Yun et al., 2021). Our rates are similar for both with- and without-replacement orderings (up to small constants), meaning they do not indicate a clear advantage for either. However, we believe they are far from tight, leaving interesting open questions in this direction.
@@
The NLMS algorithm is a classical adaptive filtering method. In its simplest version (Slock, 1993), the method perfectly fits a single—usually noisy—random sample at a time, using the same update rule as the Kaczmarz method (and thus, as our continual Scheme 1 in a rank-$1$ case). There also exists a more complex version of this method, which uses more samples per update (Sankaran and Beex, 2000). Both papers give strong $\mathcal{O}(1/k)$ MSE rates in the noiseless setting (matching our realizable setting). However, they assume a very limited data model, where the sampled vectors are either orthogonal or identical up-to-scaling. Under such conditions, Evron et al. (2022) showed that there is no forgetting (of previously learned tasks), implying that the MSE decays as the number of tasks still unseen at time $k$.

### v2

@@ 6.1 A Universal $\mathcal{O}(1/\sqrt[4]{k})$ Rate for General Projections Onto Convex Sets
Projections Onto Convex Sets (POCS) is a classical method that iteratively projects onto closed convex sets to find a point in their intersection [22, 7]. Formally,
@@
       $\mathbf{w}_{t}\leftarrow\bm{\Pi}_{\tau(t)}(\mathbf{w}_{t-1})\triangleq\operatorname*{argmin}_{\mathbf{w}\in\mathcal{C}_{\tau(t)}}\left\|{\mathbf{w}-\mathbf{w}_{t-1}}\right\|$
@@
Consider $T$ arbitrary (nonempty) closed convex sets $\mathcal{C}_{1},\dots,\mathcal{C}_{T}$, initial point $\mathbf{w}_{0}\in\mathbb{R}^{d}$, and ordering $\tau$. Define $f_{m}(\mathbf{w})=\frac{1}{2}\left\|{\mathbf{w}-\bm{\Pi}_{m}(\mathbf{w})}\right\|^{2},\forall m\in\left[{T}\right]$. Then,
@@
    The POCS update is equivalent to an SGD step: $\mathbf{w}_{t}=\bm{\Pi}_{\tau(t)}(\mathbf{w}_{t-1})=\mathbf{w}_{t-1}-\nabla_{\mathbf{w}}f_{\tau(t)}(\mathbf{w}_{t-1})$.
@@
Theorem 6.1 (Universal POCS Rate).
Consider the conditions of Reduction 3 and assume a nonempty intersection ${\mathcal{C}_{\star}=\bigcap_{m=1}^{T}\mathcal{C}_{m}\neq\varnothing}$. Then, under a random ordering with or without replacement, the expected “residual” of Scheme 4 after $\forall k\geq 1$ iterations ($k\in\left[{T}\right]$ without replacement) is bounded as,
$\mathbb{E}_{\tau}\bigg{[}\frac{1}{2T}\sum_{m=1}^{T}\left\|{\mathbf{w}_{k}-\bm{\Pi}_{m}(\mathbf{w}_{k})}\right\|^{2}\bigg{]}=\mathbb{E}_{\tau}\bigg{[}\frac{1}{2T}\sum_{m=1}^{T}\mathrm{dist}^{2}(\mathbf{w}_{k},\mathcal{C}_{m})\bigg{]}\leq\frac{7}{\sqrt[4]{k}}\min_{\mathbf{w}\in\mathcal{C}_{\star}}\left\|{\mathbf{w}_{0}-\mathbf{w}}\right\|^{2}\,.$
To the best of our knowledge, this is the first universal rate in the POCS literature, independent of problem parameters such as its regularity or complexity, as demonstrated next in Section 6.2. Universal rates are only achievable when analyzing individual distances, i.e., $f_{m}(w)=\mathrm{dist}^{2}(\mathbf{w},\mathcal{C}_{m})=\left\|{\mathbf{w}-\bm{\Pi}_{m}(\mathbf{w})}\right\|^{2}$, rather than the distance to the intersection, i.e., $\mathrm{dist}^{2}(\mathbf{w},\mathcal{C}_{\star})$. In machine learning, the squared distance from individual sets is linked to important losses like MSE in regression or squared hinge loss in classification [15, 16], naturally leading to our next continual model.
@@ 6.2 A Universal $\mathcal{O}(1/\sqrt[4]{k})$ Rate for Random Orderings in Continual Linear Classification
Regularization methods are commonly used to prevent forgetting in CL [see 33]. Evron et al. [16] studied a regularized linear model for continual classification. They considered $T\geq 2$ jointly separable, binary classification tasks, defined by datasets $S_{1},\dots,S_{T}$ consisting of vectors ${\mathbf{x}\in\mathbb{R}^{d}}$ and their labels $y\in\left\{-1,+1\right\}$. They proved that a weakly-regularized scheme implicitly applies sequential max-margin projections. That is, in the limit as $\lambda\to 0$, the iterates of the two following schemes align in direction, enabling the study of continual classification through projection algorithms.
@@
   For each iteration $t=1,\dots,k$:
@@
   For each iteration $t=1,\dots,k$:
       $\mathbf{w}_{t}\leftarrow\bm{\Pi}_{\tau(t)}(\mathbf{w}_{t-1})\triangleq\operatorname*{argmin}_{\mathbf{w}\in\mathcal{C}_{\tau(t)}}\left\|{\mathbf{w}-\mathbf{w}_{t-1}}\right\|$ where $\mathcal{C}_{m}\!\triangleq\!\left\{\mathbf{w}\!\in\!\mathbb{R}^{d}\mid y\mathbf{w}^{\top}\mathbf{x}\geq 1,~{}\forall(\mathbf{x},y)\!\in\!S_{m}\right\}$
They studied forgetting under several orderings, using an equivalent of our Definition 2.4:
$F_{\tau}(k)=\frac{1}{k}\sum_{t=1}^{k}\big{(}{\mathcal{L}_{\tau(t)}(\mathbf{w}_{k})-\mathcal{L}_{\tau(t)}(\mathbf{w}_{t})}\big{)}\leq\frac{R^{2}}{2k}\sum_{t=1}^{k}\left\|{\mathbf{w}_{k-1}-\bm{\Pi}_{\tau(t)}(\mathbf{w}_{k-1})}\right\|^{2}\,.$
The combination of our POCS rate (Theorem 6.1) and SGD stability arguments leads to the following.
Theorem 6.2.
Under a random ordering, with or without replacement, over $T$ jointly separable tasks, the expected forgetting of the weakly-regularized Scheme 5 (at $\lambda\to 0$) after $k\geq 1$ iterations is bounded as,
$\mathbb{E}_{\tau}\big{[}F_{\tau}({k})\big{]}\leq\frac{7\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}}{\sqrt[4]{k}}\,,\quad\text{where }\mathbf{w}_{\star}\in{\operatorname*{argmin}}_{\mathbf{w}\in\mathcal{C}_{1}\cap\dots\cap\mathcal{C}_{T}}\left\|{\mathbf{w}_{0}-\mathbf{w}}\right\|^{2}\,.$
As shown in Table 2, our rate is universal while the previous one depends on $\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}$, often seen as the “complexity” of classification problems. For example, after $k=4T\left\|{\mathbf{w}_{\star}}\right\|^{2}R^{2}$ iterations, the existing (normalized) rate is $e^{-1}$, while ours is potentially much smaller: $\frac{5}{T^{1/4}\sqrt{\left\|\mathbf{w}_{\star}\right\|R}}$.
@@
Evron et al. [16] $\exp\Big{(}{-\frac{k}{4T\|{\mathbf{w}_{\star}}\|^{2}R^{2}}}\Big{)}$ $\frac{T^{2}}{\sqrt{k}}\wedge\exp\Big{(}{-\frac{k}{16T^{2}\|{\mathbf{w}_{\star}}\|^{2}R^{2}}}\Big{)}$
@@ 7 Discussion
Our work established a reduction from continual linear regression to the (block) Kaczmarz method and then to “stepwise-optimal” SGD. This enabled the development of analytic tools for last-iterate SGD schemes, leading to significantly improved and even universal rates for random orderings in continual learning and the Kaczmarz method. Our main results are summarized in Tables 1 and 3.
Much of the related work has been covered throughout the paper. A further discussion of related work can be found in App. A. Here, we briefly highlight additional aspects of our work.
@@
Many popular continual benchmarks in deep learning implicitly assume a random ordering, such as the permuted MNIST benchmark [33]. Our paper shows that in sufficiently long task sequences, random ordering is enough to prevent catastrophic forgetting, and the training loss goes to zero, even in the worst case. In accordance with our results, Lesort et al. [39] examined a random CL benchmark—in which a subset of classes is randomly sampled in each task—and observed that forgetting diminishes as more tasks are sampled, even while training with standard SGD (without any modifications to mitigate forgetting). This suggests that random orderings may contaminate continual learning benchmarks, making it harder to isolate the algorithmic effects being tested. Furthermore, real-world tasks often change gradually, not adhering to random orderings. Such “gradually evolving” datasets might be more challenging and relevant as continual benchmarks.
@@
In Section 3.1 we revisited known connections between continual regression and the Kaczmarz method [15]. We broadened this connection in Section 3.2, bridging the block Kaczmarz method and “stepwise-optimal” SGD, thus applying our novel SGD bounds to the Kaczmarz method. Using Kaczmarz terminology, given a system $\mathbf{A}\mathbf{x}=\mathbf{b}$ consisting of $T$ blocks of an average rank $\bar{r}$ where $\mathbf{A}_{m}\in\mathbb{R}^{{n_{m}}\times d},\mathbf{b}_{m}\in\mathbb{R}^{n_{m}}$, our rates from Section 4 can be summarized as $\mathbb{E}_{\tau}\big{[}\frac{1}{2T}\sum_{m=1}^{T}\big{\|}{\mathbf{A}_{m}\mathbf{x}_{k}-\mathbf{b}_{m}}\big{\|}^{2}\big{]}=\mathcal{O}\big{(}{\min\big{(}{k^{-1/4},\frac{1}{k}\sqrt{d-\bar{r}},\frac{1}{k}\sqrt{T\bar{r}}}\big{)}}\big{)}$ for random orderings with replacement and $\mathcal{O}\big{(}{\min\big{(}{k^{-1/4},\frac{1}{k}\left({d-\bar{r}}\right)}\big{)}}\big{)}$ without replacement. Note that we bounded the loss, rather than the “error” $\left\|{\mathbf{w}_{k}-\mathbf{w}_{\star}}\right\|^{2}$, thus enabling the derivation of rates independent of quantities like the condition number that can make convergence arbitrarily slow.
@@
The seminal work of Strohmer and Vershynin [66] proposed a Kaczmarz variant that samples rows with probability proportional to their squared norm. Our approach accommodates non-uniform sampling, including norm based ones. This may tighten Theorem 4.2, replacing the dependence on the maximum row norm $R$ with the average one. In the block version, both uniform and non-uniform variants exist [50, 21]. There also, our approach should yield better bounds when weighting blocks by to their norms, alleviating the dependence on the maximum norm.
@@
We narrowed the gap between existing lower and upper worst-case bounds for random orderings in continual linear regression (see Table 1). However, a considerable gap remains between $\Omega(1/k)$ and $\mathcal{O}({1/k^{1/4}})$. Generally, we conjecture that the last-iterate SGD rates can be improved beyond those in Theorem 5.1, and that Theorem E.4 can be extended to the multi-epoch setup. Following our reductions (Section 3), improved rates for “stepwise-optimal” SGD rates would immediately refine the bounds for continual linear regression and classification.
Acknowledgments and Disclosure of Funding
@@ Appendix A Related Work
For the general (non-realizable) smooth stochastic setup, the recent work of Liu and Zhou [43] was the first (and only, to our knowledge) to provide upper bounds on the convergence rate of the last SGD iterate. While their bounds are applicable in the realizable setting, they require non-constant step sizes to obtain non-trivial convergence, and are therefore not useful for our purposes (see Table 3). Our analysis technique in Section 5 borrows from the work of Shamir and Zhang [62, also mentioned in Table 3] which, in fact, belongs to the comparatively-richer line of work on the non-smooth setting [62, 30, 71, 43]. Notably, SGD in a stochastic non-realizable (either smooth or non-smooth) setup requires uniformly bounded noise assumptions, and generally cannot accommodate a constant step size independent of the optimization horizon.
Table 3: State-of-the-art Loss Bounds for Fixed-Step-Size SGD. We consider stochastic convex optimization with an objective $\bar{f}(\mathbf{w})\!\triangleq\!\mathbb{E}_{\xi}f(\mathbf{w};\xi)$, where $f(\cdot;\xi)$ is $\beta$-smooth almost surely, $\sigma^{2}\geq\mathbb{E}\|\nabla f(\mathbf{w};\xi)-\nabla\bar{f}(\mathbf{w})\|^{2}$, $\sigma_{\star}^{2}\triangleq\mathbb{E}\|\nabla f(\mathbf{w}_{\star};\xi)-\nabla\bar{f}(\mathbf{w}_{\star})\|^{2}$, and $G>0$ is such that $\|\nabla f(\mathbf{w};\xi)\|\leq G$ for any $\mathbf{w}$ and $\xi$.
Dependence on constant numerical factors and the distance to an optimal solution is suppressed.
Stochastic (*) Shamir and Zhang [62] $\frac{1}{\eta T}+\eta G^{2}\log T$
Deterministic Smooth ($\sigma=0$) Nesterov [52] $\frac{1}{(2-\eta\beta)\eta T}$
Stochastic Smooth Lan [37] $\frac{1}{\eta T}+\eta\sigma^{2}$
Liu and Zhou [43] $\frac{1}{\eta T}+\eta\sigma^{2}\log T$
Stochastic Smooth Realizable ($\sigma_{\star}=0$) Srebro et al. [65] $\frac{1}{(1-\eta\beta)\eta T}$
Stochastic Regression Realizable ($\sigma_{\star}=0$) Bach and Moulines [1] $\frac{1}{\eta T}$
Varre et al. [67] $\frac{1}{(1-2\eta\beta\log T)\eta T}$
Ours (2025)  $\frac{1}{(2-\eta\beta)\eta T^{1-\eta\beta(1-\eta\beta/4)}}$
(*) They consider bounded domains [62]; Orabona [53], Liu and Zhou [43] obtain similar bounds for the unconstrained case. For non-fixed step sizes, Jain et al. [30] obtain minimax optimal bounds without log factors.
@@
Our analysis for SGD without-replacement is related to a long line of work primarily focused on the average iterate convergence rates [e.g., 58, 48, 59, 56, 47, 11, 10]. For the non-strongly convex case, near-optimal bounds (for the average iterate) have been established for the general smooth case [48, 47]. In a subsequent work, Cai et al. [10] refined the dependence on problem parameters for the smooth realizable case (among others). Guarantees for the last iterate have only been established recently by Liu and Zhou [44] and Cai and Diakonikolas [9]. However, their bounds decay with the number of epochs rather than the number of iterations and apply only to non-constant step sizes, making them inapplicable to our setting. Specifically, in a realizable $\beta$-smooth setup, after $J$ without-replacement SGD epochs over a finite sum of size $n$, Mishchenko et al. [47], Cai et al. [10] obtained an $O(\beta/J)$ bound for the average iterate with step size $\eta=1/(\beta n)$; and Liu and Zhou [44], Cai and Diakonikolas [9] derived similar bounds for the last iterate up to logarithmic factors.
@@
Another line of work related to ours studies algorithmic stability [6, 61] of gradient methods, which is the main technique we use in the proof of Theorem E.4. Our approach is similar in nature to that of Nagaraj et al. [48], Sherman et al. [63], Koren et al. [35] and primarily builds on Sherman et al. [63], who were the first to formally introduce the notion of without-replacement stability. For with-replacement SGD, [25] discussed its algorithmic stability under smooth loss functions. Later, [38], improved this bound in the realizable loss case. The case we consider—i.e., the stability of without-replacement SGD under smooth and realizable loss functions—is not covered in the existing literature.
@@
Our results in Section 4 establish universal bounds for random orderings, both with and without replacement. Both the with- and without-replacement variants converge linearly towards the minimum-norm solution $\mathbf{w}_{\star}$ [21, 24], but as we explained in Section 7, the rates can be arbitrarily slow. Recht and Ré [57] formulated a noncommutative analog of the arithmetic-geometric mean inequality that, if true, could have shown that without-replacement orderings lead to faster loss convergence than with-replacement orderings in Kaczmarz methods, and consequently in continual linear regression. Years later, Lai and Lim [36] proved that this inequality does not hold in general [see also 12]. Moreover, as in other areas, empirical studies found that row shuffling followed by cyclic orderings performs as well as i.i.d. orderings [54]. This naturally connects to interesting observations and open questions regarding various forms of shuffled SGD [5, 70]. Our rates are similar for both with- and without-replacement orderings (up to small constants), meaning they do not indicate a clear advantage for either. However, we believe they are far from tight, leaving interesting open questions in this direction.
@@
The NLMS algorithm is a classical adaptive filtering method. In its simplest version [64], the method perfectly fits a single—usually noisy—random sample at a time, using the same update rule as the Kaczmarz method (and thus, as our continual Scheme 1 in a rank-$1$ case). There also exists a more complex version of this method, which uses more samples per update [60]. Both papers give strong $\mathcal{O}(1/k)$ MSE rates in the noiseless setting (matching our realizable setting). However, they assume a very limited data model, where the sampled vectors are either orthogonal or identical up-to-scaling. Under such conditions, Evron et al. [15] showed that there is no forgetting (of previously learned tasks), implying that the MSE decays as the number of tasks still unseen at time $k$.